How to graph 4x + 4y = 12
Answer by mananth(16946)

4x+4y=12
4y=-4x+12
/4
y= -x+3
find the Y & X intercepts. Plot the points and draw the line .
when x= 0 y=3, (0,3)
when y = 0 x=3, (3,0)